Three extractions with 50.0 mL portions of chloroform removed 97% of the solute from 200 mL aqueous solution, Calculate the distribution coefficient of the solute. Calculate the no of extractions required to extract 90% of the solute with 50.0 ml portions of chloroform
1
Expert's answer
2016-09-22T11:48:03-0400
The relation between the quantity removed from an aqua solution and the distribution coefficient follows the equation Rn = 1 - (1 + Dr)^(-n). r is a relation of volumes Vaq/Vorg = 200 ml/50 ml = 4. Thus, D can be divided from the main formula as D = ((1/(1 - R3))^1/3 - 1)/r = ((1/0.03)^1/3 - 1)/4 = 0.55. On the other hand, the number of portions is n = lg(1/(1 - Rn) - 1 - Dr) = lg(1/(1 - 0.97) - 1 - 0.55*4) = 2.98 or practically 3.